OpenStudy (anonymous):

Simplify: a^7 b^8 /a^4b^4 Would the answer be a11/b12? For some reason, I had the idea that you were always supposed to add the exponents. But since this is division would you subtract the exponents instead?

OpenStudy (callisto):

\[\frac{x^a}{x^b}=x^{a-b}\]

OpenStudy (callisto):

Reason: \[x^a \cdot x^b = x^{a+b}\]\[\frac{1}{x^b} = x^{-b}\] So, \[\frac{x^a}{x^b} = x^a \frac{1}{x^b} = x^{a} \cdot x^{-b}=x^{a+(-b)}=x^{a-b}\]

OpenStudy (callisto):

For the answers to your questions, it is not a^11 / b^12. When it is division, you will do subtraction for the power of the same base.

OpenStudy (anonymous):

Thanks for explaining it - I was trying to write it all down! :)

OpenStudy (callisto):

:O You're welcome!

OpenStudy (anonymous):

the answer is a^3b^4

OpenStudy (anonymous):

if it's a fraction you just deduct the exponent of the numerator to the exponent of the denominator..